# A Classification Benchmark for Artificial Intelligence Detection of Laryngeal Cancer from Patient Speech

### v1

@@ Abstract
Cases of laryngeal cancer are predicted to rise significantly in the coming years. Current diagnostic pathways cause many patients to be incorrectly referred to urgent suspected cancer pathways, putting undue stress on both patients and the medical system.
Artificial intelligence offers a promising solution by enabling non-invasive detection of laryngeal cancer from patient speech, which could help prioritise referrals more effectively and reduce inappropriate referrals of non-cancer patients. To realise this potential, open science is crucial. A major barrier in this field is the lack of open-source datasets and reproducible benchmarks, forcing researchers to start from scratch. Our work addresses this challenge by introducing a benchmark suite comprising 36 models trained and evaluated on open-source datasets. These models are accessible in a public repository, providing a foundation for future research. They evaluate three different algorithms and three audio feature sets, offering a comprehensive benchmarking framework. We propose standardised metrics and evaluation methodologies to ensure consistent and comparable results across future studies.
The presented models include both audio-only inputs and multimodal inputs that incorporate demographic and symptom data, enabling their application to datasets with diverse patient information. By providing these benchmarks, future researchers can evaluate their datasets, refine the models, and use them as a foundation for more advanced approaches. This work aims to provide a baseline for establishing reproducible benchmarks, enabling researchers to compare new methods against these standards and ultimately advancing the development of AI tools for detecting laryngeal cancer.
@@
Artificial Intelligence , Laryngeal Cancer , Benign Voice Disorders , Voice Classification
@@
[label1]organization=School of Computer Science, University of Leeds,city=Leeds, postcode=LS2 9JT, country=UK
@@
[label2]organization=Leeds Teaching Hospitals NHS Trust,city=Leeds, postcode=LS1 3EX, country=UK
[…]
[label3]organization=School of Mathematics, University of Leeds,city=Leeds, postcode=LS2 9JT, country=UK
@@ 1 Introduction
In 2022, over 188,000 people were diagnosed with laryngeal cancer worldwide. With cases predicted to rise by 66.2% by 2050, efficient diagnosis is necessary to reduce mortality rates (Bray et al., 2024). Early detection of laryngeal cancer improves a patient’s survival rates and increases treatment options, allowing for a better quality of life post-treatment (Cancer Research UK, 2019, 2021b). Current diagnostic procedures require invasive techniques, such as nasendoscopy or laryngoscopy, to obtain specimens for biopsy. Nasendoscopy, performed as an outpatient procedure, uses a small fibre-optic camera to view the larynx via the nose, while a direct laryngoscopy is performed under general anaesthetic in the operating theatre (NHS, 2017). These methods are uncomfortable, invasive and resource-intensive.
@@
Artificial intelligence (AI) analysis of voice has been suggested as a non-invasive screening tool for the detection of laryngeal cancer, reducing the need for invasive and uncomfortable medical tests. Such a tool may be able to screen patients with concerns regarding their voice, prioritise those at the highest risk of a cancer diagnosis, expedite their specific care pathway, and increase the accessibility of diagnosis by reducing the need for expensive medical equipment as well as reducing patient stress and the load on the medical system.
Despite prior research on detecting laryngeal cancer from patient speech, progress in this field has been limited by the scarce availability of code, models and data which are rarely made available (Wang et al., 2024; Kim et al., 2024; Chen et al., 2023). Current approaches are ad-hoc with no standardisation in model assessment, making results incomparable and inconsistent and setting back advancements in this field (Paterson et al., 2024).
This work produces a suite of 36 models resulting from the combination of three algorithms, three audio feature sets, and four combinations of input variables ($3\times 3\times 4=36$). All 36 models perform the binary classification of benign and malignant voice pathologies. All pre-trained models and code used for training have been made publicly available. We also define and justify evaluation metrics, including classification performance, inference times, and fairness testing, to standardise assessment in this domain.
@@
This work establishes a benchmark for future research, aiming to accelerate progress in developing generalizable and high-performing models for laryngeal cancer detection. Researchers can use this benchmark to evaluate new datasets, refine models, and overall improve progress in this area.
@@
The rest of this work is structured as follows: Section 2 comprehensively describes the two datasets used in this work. Section 3 explains the methods used to develop the classifiers, including audio feature extraction and algorithm choice. This section also describes the evaluation methods used in this work and suggested for future work. Section 4 presents our results and discusses the impact of different algorithms, audio features, and demographic and symptom data incorporation. Section 5 concludes this work and discusses future work.
@@ 2 Evaluative Datasets
There are a limited number of publicly available datasets in this field. In our work, we use the only two datasets that, to the best of our knowledge, are currently available for public use: the Far Eastern Memorial Hospital voice dataset (FEMH) for model training and testing and the Saarbruecken Voice Database (SVD) for external validation. This section comprehensively describes and compares these two datasets.
@@ 2.1 Far Eastern Memorial Hospital voice dataset
We split this dataset using 33% of the data for testing (635 benign, 25 malignant) and the remaining data for training (1305 benign, 35 malignant). Table 2 shows a breakdown of the samples in the testing and training sets, including the number of male and female patients and the minimum, mean, and maximum age. All recordings were taken with a sample rate of 44,100 Hz.
@@ 2.2 Saarbruecken Voice Database
The Saarbruecken Voice Database (SVD) is an open-source dataset containing over 2000 participants and over 70 different pathologies (Manfred Pützer and William, J Barry, 2007). This dataset includes recordings of patients sustaining several vowels and saying a short sentence; in this work, we use the recordings of sustained /a/ at a natural pitch and volume for consistency with the audio in the FEMH dataset. This dataset also contains the patient’s age and gender.
@@
Table 2: Demographics in the training and test sets.
2.3 Dataset comparison
It’s important to compare the two datasets used in this work to understand how potential differences may impact the classifier performance.
We first compare the age and gender of the patients in each dataset. Figure 2(a) shows the distribution of ages in the datasets. It can be seen that the patients within the set of malignant samples are older than those in the benign group. Using a Mann-Whitney U test, the age distributions in the benign and malignant classes are compared between the two datasets (0.05 significance threshold). For the benign group, there is a significant difference between the age distributions (p=5.88e-6). For the malignant group, no significant difference is found (p=0.093).
Figure 2(b) shows the percentage of male and female patients within each group for each dataset. There are many more male than female patients in the malignant group. This is to be expected since laryngeal cancer affects many more men than women on average (Cancer Research UK, 2021a). Using a Fisher-Exact test, we compare the proportions of male and female patients in the benign and malignant classes between the two datasets (0.05 significance threshold). For the benign group, there is a significant difference between the proportion of male and female patients (p=9.06e-4). For the malignant group, no significant difference is found (p=1.0).
(a) The distribution of ages in the different datasets for the benign and malignant samples.
(b) The percentages of male and female samples in the different datasets for the benign and malignant samples.
Figure 2: A comparison of patient demographics within the FEMH train and set set alongside the SVD extrenal test set.
The two datasets contain a similar frequency of malignant patients, with malignant patients making up 3.0% of the FEMH dataset and 2.8% of SVD.
@@ 3.1 Audio Feature Sets
Our benchmark incorporates three distinct audio feature sets as input to the classification models. Using raw audio signals directly is impractical due to their complexity and size, as they often contain redundant information. By performing feature extraction, we simplify the audio data, reduce dimensionality, improve training and inference efficiency, and minimise overfitting (Sharma et al., 2020; Tzanetakis, 2011). The feature sets used include acoustic features, mel frequency cepstral coefficients (MFCCs), and Wav2Vec2 feature vectors. Acoustic features and MFCCs were chosen for their widespread application in similar research (Paterson et al., 2024). Although Wav2Vec2 feature vectors have not previously been applied to throat cancer detection, they have shown effectiveness in identifying other pathologies (Wagner et al., 2023). To ensure consistency, all audio recordings were resampled to 16,000 Hz before feature extraction. This is the required sample rate for Wav2Vec2 and was applied across all feature sets.
Wav2Vec2 is a large speech model trained to perform automatic speech recognition (Baevski et al., 2020). In this work, we use Facebook’s XLSR-Wav2Vec2 model; this is a pre-trained model trained using CommonVoice and Multilingual LibriSpeech, two large multi-lingual datasets (Conneau et al., 2021; Ardila et al., 2020; Pratap et al., ). Feature vectors are extracted from the final convolutional layer of the Wav2Vec2 model (from here out referred to as FeatureStates). Since the feature embeddings have variable lengths corresponding to the varying durations of the audio recordings, we applied mean pooling to standardise the representation for each recording.
OpenSMILE is an open-source feature extraction toolkit (Eyben et al., 2010). In this study, we extract the eGeMAPSv02 feature set, which comprises 88 features, including low-level descriptors (frequency, amplitude, and spectral parameters), temporal features, and cepstral parameters (Eyben et al., 2016).
@@ 3.2 Classification
For the audio, missing feature values are imputed as the mean of the respective feature. The audio features are then scaled, and feature selection is performed using a decision tree to manage the large number of audio features across all three feature sets. This process involves building a decision tree based on the audio features and then using the calculated feature importance to select only the most relevant features for input into the classifier.
[…]
For both the audio features and the demographic and symptom data, Z-score normalisation is performed to scale the features so that they have a mean of 0 and a standard deviation of 1.
This work tests three commonly used classification algorithms: support vector machine (SVM), multilayered perceptron (MLP), and logistic regression. SVM and MLP were chosen due to their due to their established use in similar research areas (Degila et al., 2018; Miliaresi et al., 2021; Kim et al., 2024), while logistic regression was chosen for its simplicity. For each classifier, a grid search with 5-fold cross-validation was emploied during training to identify the optimal hyperparameters. The grid search explores various combinations of parameters defined in a parameter grid, training models for each combination. 5-fold cross-validation is used to evaluate the parameters by splitting the data into five subsets: the models are trained on four subsets and validated on the remaining one. This process is repeated, using a different validation subset each time. The parameter combination with the best average performance across all validation subsets is selected for the final model. Table 3 shows the parameter grids used for each algorithm.
@@
To address the dataset’s imbalance of classes, class weighting was used in the SVM and logistic regression models (King and Zeng, 2001). The Synthetic Minority Oversampling Technique (SMOTE) was used for the MLP (Chawla et al., 2002). Balanced accuracy was used as the scoring metric for training across all three algorithms.
@@ 3.3 Evaluation Methods
Evaluating a model’s performance requires a multidimensional approach. While previous research in this field has primarily focused on predictive accuracy, it is crucial to consider additional metrics, especially for models intended for medical applications. This section discusses and suggests three aspects of model performance that should be assessed: predictive performance, fairness, and prediction time.

### v2

# A Classification Benchmark for Artificial Intelligence Detection of Laryngeal Cancer from Patient Voice

@@ Abstract
Cases of laryngeal cancer are predicted to rise significantly in the coming years. Current diagnostic pathways are inefficient, putting undue stress on both patients and the medical system. Artificial intelligence offers a promising solution by enabling non-invasive detection of laryngeal cancer from patient voice, which could help prioritise referrals more effectively. A major barrier in this field is the lack of reproducible methods. Our work addresses this challenge by introducing a benchmark suite comprising 36 models trained and evaluated on open-source datasets. These models classify patients with benign and malignant voice pathologies. All models are accessible in a public repository, providing a foundation for future research. We evaluate three algorithms and three audio feature sets, including both audio-only inputs and multimodal inputs incorporating demographic and symptom data. Our best model achieves a balanced accuracy of 83.7%, sensitivity of 84.0%, specificity of 83.3%, and AUROC of 91.8%.
@@
Throat Cancer , Machine Learning , Artificial Intelligence , Speech , Vocal Pathologies
@@
[first]organization=Faculty of Engineering and Physical Sciences University of Leeds,city=Leeds, country=UK
@@
[second]organization=Ear, Nose and Throat Department Leeds Teaching Hospitals NHS Trust,city=Leeds, country=UK
@@ 1 Introduction
In 2022, over 188,000 people were diagnosed with laryngeal cancer worldwide with cases predicted to rise by 66.2% by 2050 Bray et al. (2024). Early detection of laryngeal cancer improves a patient’s survival rates and increases treatment options, allowing for a better quality of life post-treatment meaning that efficient diagnosis is essential Cancer Research UK (2019, 2021a).
Current diagnostic procedures require invasive techniques, such as nasendoscopy or laryngoscopy, to obtain specimens for biopsy. Nasendoscopy, performed as an outpatient procedure, uses a small fibre-optic camera to view the larynx via the nose, while a direct laryngoscopy is performed under general anaesthetic in the operating theatre NHS (24 Oct 2017, 9:45 a.m.). These methods are uncomfortable, invasive and resource-intensive. Reducing these would reduce discomfort for patients and save money for healthcare systems.
Current referral pipelines are also inefficient with only 2.7-4.3% of patients referred on the urgent suspected cancer pathway for head and neck cancers in the UK having a cancer diagnosis NHS Digital (2024). This means that clinicians are overburdened, and patients are caused unnecessary stress while waiting for appointments as well as having to endure uncomfortable tests. By re-prioritising low-risk patients not only would the burden on the healthcare system be eased, but the cost incurred by patients in transport to appointments and loss in income caused by taking time of work to attend appointments may be reduced.
@@
Artificial intelligence (AI) analysis of voice has been suggested as a non-invasive screening tool for the detection of laryngeal cancer, reducing the need for invasive and uncomfortable medical tests. Such a tool may be able to screen patients with concerns regarding their voice, prioritise those at the highest risk of a cancer diagnosis, expedite their specific care pathway increasing the efficiency of referral systems. This would also increase the accessibility of diagnosis by reducing the need for expensive medical equipment as well as reducing patient stress and the load on the medical system.
There have been many studies that utilise AI in the detection of throat cancer from patient voice recordings Kim et al. (2020); Miliaresi et al. (2021); Kwon et al. (2022); Wang et al. (2022); Chen et al. (2023); Paterson et al. (2023); Song et al. (2023); Za’im et al. (2023); Kim et al. (2024); Wang et al. (2024). We have previously conducted a literature review into the detection of throat cancer from patient’s voice using AI and found that, although many articles have been published in this field, current approaches are ad-hoc with no standardisation in model assessment, making results incomparable and inconsistent and setting back advancements in this field Paterson et al. (2025).
We feel that one of the most important aspects of research is producing work that is reproducible such that it can be used as a jumping point for future research, without this, research areas are prone to retreading old ground without advancing. Unfortunately, the code used to create the models are rarely made available, with the only article making their code available being our previous work Paterson et al. (2023). We also found that the datasets used in this field are often not made publicly available, meaning that the work is unable to be reproduced.
@@
This work establishes a benchmark for future research, aiming to accelerate progress in developing generalizable and high-performing models for laryngeal cancer detection. Researchers can use this benchmark to evaluate new datasets, refine models, and overall improve progress in this area. To do this, this work focuses not only on creating high performing classification models but also on producing reproducible models that may be used by future researchers. This is done by using open-source datasets, and making code available.
This work produces a suite of 36 publicly available models for the classification of malignant and benign voice pathologies. These 36 models result from the combination of three classification algorithms, three audio feature sets, and four combinations of input variables ($3\times 3\times 4=36$). All trained models and code used for training have been made publicly available. We also define and justify evaluation metrics, including classification performance, inference times, and fairness testing, to standardise assessment in this domain.
@@
The rest of this work is structured as follows: Section 2 comprehensively describes the two datasets used in this work. Section 3 explains the methods used to develop the classifiers, including audio feature extraction and algorithm choice. This section also describes the evaluation methods used in this work and suggested for future work. Section 4 presents our results and discusses the impact of different algorithms, audio features, and demographic and symptom data incorporation as well as comparing our results to those of other work in this area. This section also discusses the implementability of these models into clinical practice. Section 5 concludes this work and discusses future work.
@@ 2 Evaluative Datasets
While there are many voice datasets publicly available, very few contain cancer patients. In this work we use two publicly available datasets: the Far Eastern Memorial Hospital voice dataset (FEMH) and the Saarbruecken Voice Database (SVD). These two datasets are, to the best of our knowledge, the only two publicly available datasets that contain voice recordings of patients with laryngeal cancer.
The FEMH dataset is split into a training set used to train the classification models and a test set used to test the classification models. The SVD dataset is used for testing only. This section describes and compares these two datasets.
@@ 2.1 Far Eastern Memorial Hospital voice dataset
The Far Eastern Memorial Hospital (FEMH) dataset was produced by Wang et al. Wang et al. (2024). It contains audio recordings of 2000 individuals sustaining the vowel /a/. This dataset also contains comprehensive, structured medical records for each individual, including the sex and age of the patient, as well as 23 symptoms. All recordings have a sample rate of 44,100 Hz
@@
The FEMH dataset was used to train all of the models as well as being used for testing. We split this dataset using 33% of the data for testing (635 benign, 25 malignant) and the remaining 77% of the data was used for training (1305 benign, 35 malignant).
@@ 2.2 Saarbruecken Voice Database
The Saarbruecken Voice Database (SVD) is an open-source dataset containing over 2000 participants and over 70 different pathologies Manfred Pützer and William, J Barry (2007). This dataset includes recordings of patients sustaining several vowels and saying a short sentence; in this work, we use the recordings of sustained /a/ at a natural pitch and volume for consistency with the audio in the FEMH dataset. This dataset also contains the patient’s age and sex.
@@
2.3 Dataset comparison
It’s important to compare the two datasets used in this work to understand how differences may impact the classifier performance. Table 2 shows a breakdown of the samples in the testing and training sets for both the FEMH and SVD datasets. This table shows the number of male and female patients and the minimum, mean, and maximum age.
@@ 2.3 Dataset comparison
Table 2: Demographics in the training and test sets for both the FEMH and SVD datasets.
We first compare the age and sex of the patients in each dataset. Figure 2 shows the distribution of ages in the datasets. It can be seen that the patients in the malignant group are older than those in the benign group. Using a Mann-Whitney U test, we investigate whether the age distributions differ significantly between the FEMH and SVD datasets (0.05 significance threshold). A separate test was performed for the benign and malignant groups. For the benign group, there is a significant difference between the age distributions between the FEMH and SVD datasets (p=5.88e-6). For the malignant group, no significant difference is found (p=0.093). Therefore misclassification of benign patients in the SVD dataset may be due to this significant difference in ages, as the classifiers are trained on the FEMH dataset.
Figure 2: The distribution of ages in the different datasets for the benign and malignant samples.
Next we compare the sex of the patients in the benign and malignant group for both the FEMH and SVD datasets. Figure 3 shows the percentage of male and female patients within each group for both datasets. There are many more male than female patients in the malignant group. This is to be expected since laryngeal cancer affects many more men than women on average Cancer Research UK (2021b). Using a Fisher-Exact test, we compare the proportions of male and female patients in the benign and malignant classes between the FEMH and SVD datasets (0.05 significance threshold). For the benign group, there is a significant difference between the proportion of male and female patients in the FEMH and SVD datasets (p=9.06e-4). For the malignant group, no significant difference is found (p=1.0). Once again, this means that the misclassification of benign patients from the SVD dataset may be due to the significant difference in the proportion of male and female patients in the SVD dataset compared to the FEMH dataset.
Figure 3: The percentages of male and female samples in the different datasets for the benign and malignant samples.
The two datasets contain a similar frequency of malignant patients, with malignant patients making up 3.0% of the FEMH dataset and 2.8% of SVD. This is a significant data imbalance which is common within this area of research Kwon et al. (2022); Wang et al. (2022); Chen et al. (2023). This imbalance must be considered when developing and evaluating AI systems.
@@ 3 Methods
The pipeline used to create these models can be seen in Figure 4. This method outline could be used and expanded in future work. We start by extracting audio features, that are then preprocessed separately to the demographic and symptom data. After both sets of features are preprocessed, they are combined and used to train the model, in this case using grid search cross-validation. In this section we discuss the audio feature sets, preprocessing methods, and classification models used in this work. In future work however, these steps could be modified to further optimize model performance.
In this section we also suggest evaluation methods to be used in future work. These not only include predictive performance metrics but also evaluation of model fairness and prediction times which should be reported to give a better understanding of how models may be implemented into clinical practice.
@@ 3.1 Audio Feature Sets
Our benchmark incorporates three distinct audio feature sets as input to the classification models. The feature sets used are OpenSMILE acoustic features, mel frequency cepstral coefficients (MFCCs), and Wav2Vec2 feature vectors. Acoustic features and MFCCs were chosen for their widespread application in similar research Bhat and Kopparapu (2018); Grzywalski et al. (2018); Islam et al. (2018). Although Wav2Vec2 feature vectors have not previously been applied to throat cancer detection, they have shown effectiveness in identifying other pathologies Wagner et al. (2023). To ensure consistency, all audio recordings were resampled to 16,000 Hz before feature extraction. This is the required sample rate for Wav2Vec2 and was applied across all feature sets. While some work used the raw audio signals directly as input to the system we feel that this is impractical due to their complexity and size Kim et al. (2020); Wang et al. (2024). Of course, raw audio signals could be used as input to the model while using the pipeline shown in Figure 4. By performing feature extraction, we simplify the audio data, reduce dimensionality, improve training and inference efficiency, and minimise overfitting Sharma et al. (2020); Tzanetakis (2011).
Wav2Vec2 is a large speech model trained to perform automatic speech recognition Baevski et al. (2020). In this work, we use Facebook’s XLSR-Wav2Vec2 model; this is a pre-trained model trained using CommonVoice and Multilingual LibriSpeech, two large multi-lingual datasets Conneau et al. (2021); Ardila et al. (2020); Pratap et al. . Feature vectors are extracted from the final convolutional layer of the Wav2Vec2 model (from here out referred to as FeatureStates). Since the feature embeddings have variable lengths corresponding to the varying durations of the audio recordings, we applied mean pooling to standardise the representation for each recording.
OpenSMILE is an open-source feature extraction toolkit Eyben et al. (2010). In this study, we extract the eGeMAPSv02 feature set, which comprises 88 features, including low-level descriptors (frequency, amplitude, and spectral parameters), temporal features, and cepstral parameters Eyben et al. (2016).
@@
3.2 Feature Preprocessing
For the audio feature sets chosen in this work missing values are uncommon however, some acoustic features that are based on fundamental frequency in the OpenSMILE feature set (and other similar feature sets) may be missing when patients have a particularly hoarse voice. This is because when a patient’s voice is very hoarse pitch tracking algorithms can be ineffective and fail. In the case that values in the audio feature sets are missing, feature values are imputed as the mean of the respective feature. The audio features are then scaled, and feature selection is performed using a decision tree to manage the large number of audio features across all three feature sets. This process involves building a decision tree based on the audio features and then using the calculated feature importance to select only the most relevant features for input into the classifier.
[…]
For both the audio features and the demographic and symptom data, Z-score normalisation is performed to scale the features so that they have a mean of 0 and a standard deviation of 1. This is done so that features with a large scales do not overpower the classification algorithm Géron (2022).
@@ 3.3 Classification
In this work we present three commonly used classification algorithms: support vector machine (SVM), multilayered perceptron (MLP), and logistic regression. SVM and MLP were selected due to their established use in similar research areas Degila et al. (2018); Miliaresi et al. (2021); Kim et al. (2024), while logistic regression was chosen for its simplicity. Once again, in future work, any classification algorithm could be used within the classification pipeline.
While deep learning methods are commonly used within the area of AI voice classification, our literature review showed that they yielded no improvement in classification over simpler methods Paterson et al. (2025). While we did attempt more complex deep learning approaches, we did not find any that out-performed the SVM, MLP, and logistic regression in terms of classification performance. We believe that this is due to the relatively small amount of available data, should more data be made available, these methods may yield an improvement in classifier results. However, in the current data landscape, we strongly believe that the use of deep learning methods is unnecessary and likely to cause overfitting should good results be obtained.
For each classifier, a grid search with 5-fold cross-validation was employed during training to identify the optimal hyperparameters. The grid search explores various combinations of parameters defined in a parameter grid, training models for each combination. 5-fold cross-validation is used to evaluate the parameters by splitting the data into five subsets: the models are trained on four subsets and validated on the remaining one. This process is repeated, using a different validation subset each time. The parameter combination with the best average performance across all validation subsets is selected for the final model. Table 3 shows the parameter grids used for each algorithm.
@@
To address the FEMH dataset’s imbalance of classes, class weighting was used in the SVM and logistic regression models King and Zeng (2001). The Synthetic Minority Oversampling Technique (SMOTE) was used for the MLP Chawla et al. (2002). Balanced accuracy was used as the scoring metric for training across all three algorithms.
@@ 3.4 Evaluation Methods
While previous research in this field has primarily focused on predictive accuracy, it is crucial to consider additional metrics, especially for models intended for medical applications. This section discusses and suggests three aspects of model performance that should be assessed: predictive performance, fairness, and prediction time. These evaluation methods provide standardized model assessment allowing for work to be consistent and comparable.
The use of a holdout test set (i.e. a set of data not used for training but originating from the same dataset as the training data) allows for evaluation of the model’s capacity to perform on unseen data but does not allow for evaluation of the model’s generalizability to variations likely to be seen in real world implementation. Using an external test set (i.e. a set of data used for testing that originates from a different dataset to the data used in training) allows for better understanding of how the models will perform in real world settings. As such it’s important that models are evaluated on both holdout dataset (the FEMH dataset in this work) and external test sets (the SVD dataset in this work).